OpenStudy (anonymous):

I need help please, I am not sure how to approach it: sinx-cosx=0 (Find all solutions)

OpenStudy (anonymous):

Am I supposed to do the quotient identity?

OpenStudy (jdoe0001):

\(\bf sin^2(\theta)+cos^2(\theta)=1\implies sin^2(\theta)=1-cos^2(\theta) \\ \quad \\ {\color{brown}{ sin(\theta)=\pm \sqrt{1-cos^2(\theta)}}}\qquad thus \\ \quad \\ \quad \\ sin(x)-cos(x)=0\implies \sqrt{1-cos^2(x)}-cos(x)=0 \\ \quad \\ \sqrt{1-cos^2(x)}=cos(x)\implies 1-cos^2(x)=[cos(x)]^2 \\ \quad \\ 1-cos^2(x)=cos^2(x)\implies 1=2cos^2(x)\implies \cfrac{1}{2} \\ \quad \\ \pm \sqrt{\cfrac{1}{2}}=cos(x)\) then take the arcCos() to each side to get "x"

OpenStudy (jdoe0001):

hmm anyhow missing a bit \(\bf sin^2(\theta)+cos^2(\theta)=1\implies sin^2(\theta)=1-cos^2(\theta) \\ \quad \\ {\color{brown}{ sin(\theta)=\pm \sqrt{1-cos^2(\theta)}}}\qquad thus \\ \quad \\ \quad \\ sin(x)-cos(x)=0\implies \sqrt{1-cos^2(x)}-cos(x)=0 \\ \quad \\ \sqrt{1-cos^2(x)}=cos(x)\implies 1-cos^2(x)=[cos(x)]^2 \\ \quad \\ 1-cos^2(x)=cos^2(x)\implies 1=2cos^2(x)\implies \cfrac{1}{2}=cos^2(x) \\ \quad \\ \pm \sqrt{\cfrac{1}{2}}=cos(x)\)

OpenStudy (anonymous):

are you square rooting it because the sine has an exponent of 2?

OpenStudy (jdoe0001):

hm,m?

OpenStudy (jdoe0001):

what do you mean?

OpenStudy (anonymous):

I don't understand why the sine and cosine are squared